OpenStudy (anonymous):

find the derivates of this function: y=xsin(2/x)

OpenStudy (anonymous):

\[xsin(2/x)\]

OpenStudy (anonymous):

Which derivatives? The first two?

OpenStudy (anonymous):

the whole function please

OpenStudy (anonymous):

Well, how would you start? I see two components: the x and the sin (2/x). How do you derive the product of two functions?

OpenStudy (anonymous):

differentiate*

OpenStudy (anonymous):

whats the difference between derive adn differentiate?

OpenStudy (anonymous):

Nothing really; I think differentiate is the correct term.

OpenStudy (anonymous):

hint, use the product rule : (fg)'=fg'+f'g

OpenStudy (anonymous):

while we're doing these, does anyone know if i would use the same rule for g(n)=(n³-n)ln(n²+n) ?

OpenStudy (anonymous):

Sure; the easiest way I'd see is to say a(x) = (n^3 - n) and b(x) = ln(n^2 + n)

OpenStudy (anonymous):

er, swap the xs for ns there

OpenStudy (anonymous):

haha i realised that:)

OpenStudy (anonymous):

use the product rule, (fg)'=f'g+g'f and remember for the sin (2/x) you will need to use the chain rule.

OpenStudy (anonymous):

ok but whats derivitve of sin?

OpenStudy (anonymous):

d/dx sin x is cos x. Make sure to use the Chain Rule, though!